# Federated Learning With Energy Harvesting Devices: An MDP Framework

### v1

@@ III-B Problem Formulation
In this paper, we aim to optimize the device scheduling strategy $\{\beta_{n,t}\}$ and power control strategy $\{p_{n,t}\}$ to minimize the global loss $F(\boldsymbol{w}_{T+1})$ after $T$ time slots. However, directly solving this problem is challenging since the global loss is hard to express in a closed form in terms of the transmission strategy. To address this issue, we approximate the actual global loss with its convergence bound derived in Theorem 1, and formulate an optimality gap minimization problem. Moreover, we ignore the first term on the RHS of (17) since it only relates to the initial error and converges to zero as $T$ increases. The formulated problem is given by
$\underset{\{\boldsymbol{\beta}_{t},\boldsymbol{p}_{t}\}_{t=1}^{T}}{\text{min}}$ $\;\frac{\eta\rho}{2N}\sum_{t=1}^{T}A^{T-t}\sum_{n=1}^{N}\frac{1-\beta_{n,t}(1-q_{n,t}(h_{n,t},p_{n,t}))}{D_{n}}$
@@
$\underset{\{\boldsymbol{p}_{t}\}_{t=1}^{T}}{\text{min}}$ $~{}\frac{\eta\rho}{2N}\sum_{t=1}^{T}\sum_{n=1}^{N}\frac{A^{T-t}q_{n,t}(h_{n,t},p_{n,t})}{D_{n}}$ (19)
@@ IV-A MDP Formulation
    where $c_{n}({h}_{n,t},{p}_{n,t})=\frac{\eta\rho q_{n,t}(h_{n,t},p_{n,t})}{2ND_{n}}$ is the one-step cost of device $n$.
@@
$J(\pi)=\mathbb{E}_{\pi}\left[\sum_{t=1}^{T}A^{T-t}c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})|\boldsymbol{h}_{1},\boldsymbol{b}_{1}\right],$ (24)
@@
We note that the objective function in Problem 1 is a weighted sum of costs $c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})$, where the weight $A^{T-t}$ increases with time slot $t$. This formulation differs from the standard discounted MDPs, which apply a discount factor typically less than $1$ to diminish the importance of future costs over time. To address this unique inversion of discounting, we introduce a modified Bellman equation in the following subsection to compute the optimal policy for Problem 1.
@@ IV-B Optimal Policy
$=$ $\min_{\boldsymbol{p}_{t}\in\mathcal{P}_{t}}\left\{A^{T-t}c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})+\mathbb{E}\left[V_{t+1}(\boldsymbol{h}_{t+1},\boldsymbol{b}_{t+1})|\boldsymbol{h}_{t},\boldsymbol{b}_{t},\boldsymbol{p}_{t}\right]\right\},$
@@
$V_{T}(\boldsymbol{h}_{T},\boldsymbol{b}_{T})=$ $\min_{\boldsymbol{p}_{T}\in\mathcal{P}_{T}}\left\{c(\boldsymbol{h}_{T},\boldsymbol{p}_{T})\right\},$ (27)
@@
$\boldsymbol{p}^{*}_{t}\!=$ $\operatorname*{arg\,min}_{\boldsymbol{p}_{t}\in\mathcal{P}_{t}}\!\left\{\!A^{T-t}c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})\!+\!\mathbb{E}\!\left[V_{t+1}\!(\boldsymbol{h}_{t+1},\boldsymbol{b}_{t+1})|\boldsymbol{h}_{t},\boldsymbol{b}_{t},\boldsymbol{p}_{t}\right]\right\}\!\!,$ (28)
@@
$\boldsymbol{p}^{*}_{T}=$ $\operatorname*{arg\,min}_{\boldsymbol{p}_{T}\in\mathcal{P}_{T}}\left\{c(\boldsymbol{h}_{T},\boldsymbol{p}_{T})\right\}.$ (29)
@@
$=A^{T-t}c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})+\mathbb{E}\left[V_{t+1}(\boldsymbol{h}_{t+1},\boldsymbol{b}_{t+1})|\boldsymbol{h}_{t},\boldsymbol{b}_{t},\boldsymbol{p}_{t}\right].$
@@
It is evident that the first term of $Z_{n,t}(b_{n,t},p_{n,t})$, i.e., $A^{T-t}c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})$, is submodular in $(b_{n,t},p_{n,t})$ since it is independent of $b_{n,t}$. Then, define $\boldsymbol{x}_{t}=\boldsymbol{b}_{t}-\boldsymbol{e}_{t}$ with its $n$-th entry $x_{n,t}={b}_{n,t}-{e}_{n,t}$. Given fixed $\left\{b_{m,t},m\in[N]\backslash n\right\}$, we denote the second term of $Z_{n,t}(b_{n,t},p_{n,t})$ as
@@
Now let $x_{n,t}=b_{n,t}^{-}-e_{n,t}^{+}$, $x_{n,t}^{\prime}=b_{n,t}^{-}-e_{n,t}^{-}$, and $\epsilon=b_{n,t}^{+}-b_{n,t}^{-}$, where $e_{n,t}^{+}$ and $e_{n,t}^{-}$ correspond to the energy consunption under the transmisson power $p_{n,t}^{+}$ and $p_{n,t}^{-}$, respectively. Then, we can obtain that the second term of $Z_{n,t}(b_{n,t},p_{n,t})$ is also submodular in $(b_{n,t},p_{n,t})$. Thus, $p_{n,t}^{*}$ is non-decreasing in $b_{n,t}$. ∎
@@
$V_{t}(\boldsymbol{h}_{t},\boldsymbol{b}_{t})\!=\!A^{T-t}c(\boldsymbol{h}_{t},\boldsymbol{p}^{\#}_{t})\!+\!\mathbb{E}\!\left[\!V_{t+1}(\boldsymbol{h}_{t+1},\boldsymbol{b}_{t+1})|\boldsymbol{h}_{t},\boldsymbol{b}_{t},\boldsymbol{p}^{\#}_{t}\!\right]\!\!,$ (36)
@@ V-A Device Level Decomposition
$=$ $\mathbb{E}_{\pi}\left[\sum_{t=1}^{T}\left(A^{T-t}c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})+\lambda\left(\sum_{n=1}^{N}\mathbbm{1}_{\left\{p_{n,t}>0\right\}}-R\right)\right)\right].$
@@
Note that the dimension of the state space $\mathcal{S}$ is finite and the one-step cost is bounded below, i.e., $c(\boldsymbol{h},\boldsymbol{p})\geq 0,\forall\boldsymbol{h},\boldsymbol{p}$. Under these conditions, there exists an optimal Lagrange multiplier $\lambda^{*}$ such that $J(\overline{\pi}^{*})=\mathcal{L}^{*}(\lambda^{*})$ and $\lambda^{*}\left(R(\overline{\pi}_{\lambda^{*}})-R\right)=0$ [31, Theorem 12.8]. Thus, the optimal relaxed policy $\overline{\pi}^{*}$ can be computed by using a two-stage iterative algorithm: 1) For a given $\lambda$, we find the $\lambda$-optimal policy; 2) we update $\lambda$ by applying a bisection method. These two steps are detailed in the following.
@@ V-A1 Optimal Policy for a Fixed Lagrange Multiplier
$=$ $\mathbb{E}_{\pi_{n}}\left[\sum_{t=1}^{T}\left(A^{T-t}c_{n}({h}_{n,t},{p}_{n,t})+\lambda\mathbbm{1}_{\left\{p_{n,t}>0\right\}}-\frac{\lambda R}{N}\right)\right],$
@@ V-A2 Determination of the Optimal Lagrange Multiplier
According to [32, Lemma 3.1], $R(\overline{\pi}_{\lambda})$ decreases as $\lambda$ increases. Intuitively, the Lagrange multiplier $\lambda$ can be interpreted as the cost of scheduling one device; thus, increasing $\lambda$ imposes a greater penalty on the scheduling cost in the Lagrangian. Our goal is to identify the smallest value of $\lambda$ that satisfies the relaxed bandwidth constraint $\overline{\textup{C3}}$. Formally, the optimal Lagrange multiplier is defined as
@@ 12            Set $\lambda^{+}=\lambda$;
Remark 2.
V-B Truncation Policy With Hard Bandwidth Constraint
@@ V-C Asymptotic Optimality of the Relax-and-Truncate Policy
$J(\tilde{\pi})-J(\pi^{*})\leq\frac{\eta\rho T}{2\kappa D_{\textup{min}}\sqrt{N}}=\mathcal{O}\left(\frac{1}{\sqrt{N}}\right),$ (44)
where $D_{\textup{min}}=\min_{n}\left\{D_{n}\right\}$.
@@ VI Structure-Enhanced Deep Reinforcement Learning
In this section, we address a more general scenario in which the statistical knowledge of channels and harvested energy is unknown. To this end, we develop a novel structure-enhanced Deep Q-Network (SE-DQN) algorithm that leverages the monotone structure of the optimal policy to improve offline training performance.
@@ VI-A Deep Reinforcement Learning
$Q(\boldsymbol{h}_{t},\boldsymbol{b}_{t},\boldsymbol{p}_{t})$ (45)
$=$ $c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})\!+\!A^{\!-1}\mathbb{E}\!\!\left[\min_{\!\boldsymbol{p}_{t+1}\!\in\mathcal{P}_{t+1}}\!\!\!\left\{\!Q(\boldsymbol{h}_{t+1},\boldsymbol{b}_{t+1},\boldsymbol{p}_{t+1})\!\right\}\!\!|\boldsymbol{h}_{t},\boldsymbol{b}_{t},\boldsymbol{p}_{t}\right]\!\!.$
@@
$y_{t}=c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})+A^{-1}\min_{\!\!\boldsymbol{p}_{t+1\!}\in\mathcal{P}_{t+1}\!\!}Q(\boldsymbol{h}_{t+1},\boldsymbol{b}_{t+1},\boldsymbol{p}_{t+1};\boldsymbol{\theta}^{\prime}),$ (47)
@@ VII-A1 Learning Model Setting
where $C$ represents the total number of label categories in the dataset, $\boldsymbol{w}_{c}$ denotes the model parameter vector associated with label category $c\in[C]$, and $\alpha_{r}=0.01$. For the CIFAR-10 dataset, we train a convolutional neural network (CNN) with the following structure: four $3\times 3$ convolution layers each with 64 channels and followed by a $2\times 2$ max pooling layer, a fully connected layer with 120 units, and finally a 10-unit softmax output layer. Each convolution or fully connected layer is activated by a ReLU function. Both the MNIST and CIFAR-10 datasets are randomly and uniformly partitioned into $N$ disjoint subsets, and each device holds one subset as its local training dataset. We set the learning rate $\eta$ as 0.01 for the MNIST dataset and 0.05 for the CIFAR-10 dataset.
@@ VII-D Comparison With Benchmark Schemes
    $\underset{p_{n,t}}{\text{min}}{}{}$ $\frac{V\eta\rho A^{T-t}q_{n,t}(h_{n,t},p_{n,t})}{2ND_{n}}+Q_{n,t}e_{n,t}$ (55)
@@
    where $V>0$ is a tradeoff hyperparameter, and $Q_{n,t}$ is the virtual queue of device $n$. The queuing dynamics of $Q_{n,t}$ is given by
    $Q_{n,t+1}=Q_{n,t}+e_{n,t}-u_{n,t}.$ (56)
In Fig. 5, we compare the learning performance of the proposed algorithms with benchmark schemes on the MNIST dataset. We consider $N=200$ edge devices with a total bandwidth of $R=100$. As observed, the proposed algorithms achieve near-optimal learning performance and outperform other benchmarks. The greedy optimality gap minimization scheme myopically uses up all remaining energy at each time slot, leading to poor performance in subsequent time slots. The Lyapunov-based online optimization scheme does not utilize the statistical knowledge of system dynamics, e.g., the distributions of the channel gains and the amount of harvested energy, resulting in poor long-term performance. In contrast, the proposed algorithms leverage the previously known or learned knowledge of channel and energy dynamics to design the transmission policy from a long-term perspective, thereby improving the learning performance.
Fig. 6 shows a similar comparison on the CIFAR-10 dataset. To demonstrate the wide applicability of the proposed algorithms, we implement CNNs as the classifier model, which does not satisfy Assumptions 1 and 2. Despite this, similar empirical performance trends can be observed and our proposed algorithms still outperform other benchmarks, except for the optimal benchmark scheme.
Fig. 5: Test accuracy versus the number of time slots under different transmission strategies on the MNIST dataset.

### v2

@@ III-B Problem Formulation
In this paper, we aim to optimize the device scheduling strategy $\{\beta_{n,t}\}$ and power control strategy $\{p_{n,t}\}$ to minimize the expected global loss after $T$ time slots. However, directly solving this problem is challenging since the global loss is hard to express in a closed form in terms of the transmission strategy. To address this issue, we approximate the actual global loss with its convergence bound derived in Theorem 1, and formulate an optimality gap minimization problem. In particular, we ignore the first four terms on the RHS of (LABEL:convergence_bound), as they are independent of the device scheduling and transmission power and go to zero as $T$ increases. Hence, the resulting optimization problem is formulated as follows:
$\underset{\{\boldsymbol{\beta}_{t},\boldsymbol{p}_{t}\}_{t=1}^{T}}{\text{min}}$ $\;\sum_{t=1}^{T}\sum_{n=1}^{N}\frac{2LG^{2}\left(1-\beta_{n,t}(1-q_{n,t}(h_{n,t},p_{n,t}))\right)}{N\sqrt{KT}}$
@@
$\underset{\{\boldsymbol{p}_{t}\}_{t=1}^{T}}{\text{min}}$ $~{}\sum_{t=1}^{T}\sum_{n=1}^{N}\frac{2LG^{2}q_{n,t}(h_{n,t},p_{n,t})}{N\sqrt{KT}}$ (21)
@@ IV-A MDP Formulation
    where $c_{n}({h}_{n,t},{p}_{n,t})=\frac{2LG^{2}q_{n,t}(h_{n,t},p_{n,t})}{N\sqrt{KT}}$ is the one-step cost of device $n$.
@@
$J(\pi)=\mathbb{E}_{\pi}\left[\sum_{t=1}^{T}c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})|\boldsymbol{h}_{1},\boldsymbol{b}_{1}\right],$ (26)
@@
With Problem 1 formally established as a finite-horizon MDP, we next characterize the optimal transmission policy by leveraging dynamic programming. Specifically, in the following subsection, we derive the optimal transmission policy via the Bellman equation and analyze its structural properties.
@@ IV-B Optimal Policy
$=$ $\min_{\boldsymbol{p}_{t}\in\mathcal{P}_{t}}\left\{c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})+\mathbb{E}\left[V_{t+1}(\boldsymbol{h}_{t+1},\boldsymbol{b}_{t+1})|\boldsymbol{h}_{t},\boldsymbol{b}_{t},\boldsymbol{p}_{t}\right]\right\},$
@@
$V_{T}(\boldsymbol{h}_{T},\boldsymbol{b}_{T})=$ $\min_{\boldsymbol{p}_{T}\in\mathcal{P}_{T}}c(\boldsymbol{h}_{T},\boldsymbol{p}_{T}),$ (29)
@@
$\boldsymbol{p}^{*}_{t}\!=$ $\operatorname*{arg\,min}_{\boldsymbol{p}_{t}\in\mathcal{P}_{t}}\!\left\{c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})\!+\!\mathbb{E}\!\left[V_{t+1}\!(\boldsymbol{h}_{t+1},\boldsymbol{b}_{t+1})|\boldsymbol{h}_{t},\boldsymbol{b}_{t},\boldsymbol{p}_{t}\right]\right\}\!\!,$ (30)
@@
$\boldsymbol{p}^{*}_{T}=\operatorname*{arg\,min}_{\boldsymbol{p}_{T}\in\mathcal{P}_{T}}\;c(\boldsymbol{h}_{T},\boldsymbol{p}_{T}).$ (31)
@@
$=c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})+\mathbb{E}\left[V_{t+1}(\boldsymbol{h}_{t+1},\boldsymbol{b}_{t+1})|\boldsymbol{h}_{t},\boldsymbol{b}_{t},\boldsymbol{p}_{t}\right].$
@@
It is evident that the first term of $Z_{n,t}(b_{n,t},p_{n,t})$, i.e., $c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})$, is submodular in $(b_{n,t},p_{n,t})$ since it is independent of $b_{n,t}$. Then, define $\boldsymbol{x}_{t}=\boldsymbol{b}_{t}-\boldsymbol{e}_{t}$ with its $n$-th entry $x_{n,t}={b}_{n,t}-{e}_{n,t}$. Given fixed $\left\{b_{m,t},m\in[N]\backslash n\right\}$, we denote the second term of $Z_{n,t}(b_{n,t},p_{n,t})$ as
@@
Now let $x_{n,t}=b_{n,t}^{-}-e_{n,t}^{+}$, $x_{n,t}^{\prime}=b_{n,t}^{-}-e_{n,t}^{-}$, and $\epsilon=b_{n,t}^{+}-b_{n,t}^{-}$, where $e_{n,t}^{+}$ and $e_{n,t}^{-}$ correspond to the energy consumption under the transmisson power $p_{n,t}^{+}$ and $p_{n,t}^{-}$, respectively. Then, we can obtain that the second term of $Z_{n,t}(b_{n,t},p_{n,t})$ is also submodular in $(b_{n,t},p_{n,t})$. Thus, $p_{n,t}^{*}$ is non-decreasing in $b_{n,t}$. ∎
@@
$V_{t}(\boldsymbol{h}_{t},\boldsymbol{b}_{t})=c(\boldsymbol{h}_{t},\boldsymbol{p}^{\#}_{t})+\mathbb{E}\!\left[V_{t+1}(\boldsymbol{h}_{t+1},\boldsymbol{b}_{t+1})|\boldsymbol{h}_{t},\boldsymbol{b}_{t},\boldsymbol{p}^{\#}_{t}\right],$ (38)
@@ V-A Device Level Decomposition
$=$ $\mathbb{E}_{\pi}\left[\sum_{t=1}^{T}\left(c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})+\lambda\left(\sum_{n=1}^{N}\mathbbm{1}_{\left\{p_{n,t}>0\right\}}-R\right)\right)\right].$
@@
Note that the dimension of the state space $\mathcal{S}$ is finite and the one-step cost is bounded below, i.e., $c(\boldsymbol{h},\boldsymbol{p})\geq 0,\forall\boldsymbol{h},\boldsymbol{p}$. Under these conditions, there exists an optimal Lagrange multiplier $\lambda^{*}$ such that $J(\overline{\pi}^{*})=\mathcal{L}^{*}(\lambda^{*})$ and $\lambda^{*}\left(R(\overline{\pi}_{\lambda^{*}})-R\right)=0$ [33, Theorem 12.8]. Thus, the optimal relaxed policy $\overline{\pi}^{*}$ can be computed by using a two-stage iterative algorithm: 1) For a given $\lambda$, we find the $\lambda$-optimal policy; 2) we update $\lambda$ by applying a bisection method. These two steps are detailed in the following.
@@ V-A1 Optimal Policy for a Fixed Lagrange Multiplier
$=$ $\mathbb{E}_{\pi_{n}\!\!}\left[\sum_{t=1}^{T}\left(c_{n}({h}_{n,t},{p}_{n,t})\!+\!\lambda\mathbbm{1}_{\left\{p_{n,t}>0\right\}}\!-\!\frac{\lambda R}{N}\right)\!\right]\!,$
@@ V-A2 Determination of the Optimal Lagrange Multiplier
According to [34, Lemma 3.1], $R(\overline{\pi}_{\lambda})$ decreases as $\lambda$ increases. Intuitively, the Lagrange multiplier $\lambda$ can be interpreted as the cost of scheduling one device; thus, increasing $\lambda$ imposes a greater penalty on the scheduling cost in the Lagrangian. Our goal is to identify the smallest value of $\lambda$ that satisfies the relaxed bandwidth constraint $\overline{\textup{C3}}$. Formally, the optimal Lagrange multiplier is defined as
@@
Remark 2.
V-B Truncation Policy With Hard Bandwidth Constraint
@@ V-C Asymptotic Optimality of the Relax-and-Truncate Policy
$J(\tilde{\pi})-J(\pi^{*})\leq\frac{2LG^{2}\sqrt{T}}{\kappa\sqrt{NK}}=\mathcal{O}\left(\frac{1}{\sqrt{N}}\right).$ (46)
@@ VI Structure-Enhanced Deep Reinforcement Learning
In this section, we address a more general scenario in which the statistical knowledge of channels and harvested energy is unknown. To this end, we develop a structure-enhanced Deep Q-Network (SE-DQN) algorithm that leverages the monotone structure of the optimal policy to improve offline training performance.
@@ VI-A Deep Reinforcement Learning
$Q(\boldsymbol{h}_{t},\boldsymbol{b}_{t},\boldsymbol{p}_{t})=c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})+\min_{\boldsymbol{p}_{t+1}\in\mathcal{P}_{t+1}}Q(\boldsymbol{h}_{t+1},\boldsymbol{b}_{t+1},\boldsymbol{p}_{t+1}).$ (47)
@@
$y_{t}=c(\boldsymbol{h}_{t},\boldsymbol{p}_{t})+\min_{\boldsymbol{p}_{t+1}\in\mathcal{P}_{t+1}}Q(\boldsymbol{h}_{t+1},\boldsymbol{b}_{t+1},\boldsymbol{p}_{t+1};\boldsymbol{\theta}^{\prime}),$ (49)
@@ VII-A1 Learning Model Setting
where $C$ represents the total number of label categories in the dataset, $\boldsymbol{w}_{c}$ denotes the model parameter vector associated with label category $c\in[C]$, and $\alpha_{r}=0.01$. For the CIFAR-10 dataset, we train a convolutional neural network (CNN) with the following structure: four $3\times 3$ convolution layers each with 64 channels and followed by a $2\times 2$ max pooling layer, a fully connected layer with 120 units, and finally a 10-unit softmax output layer. Each convolution or fully connected layer is activated by a ReLU function. Both the MNIST and CIFAR-10 datasets are divided into $N$ subsets, with each subset assigned to a distinct device as its local training dataset. We investigate both i.i.d. and non-i.i.d. data distributions. For i.i.d. scenarios, we implement uniform random data distribution across devices. To simulate realistic heterogeneity in non-i.i.d. settings, we employ a Dirichlet distribution Dir(0.8) for both the FMNIST and CIFAR-10 datasets. We set the learning rate $\eta$ as 0.01 for the MNIST dataset and 0.05 for the CIFAR-10 dataset.
@@ VII-A2 Communication Model Setting
(a) i.i.d.
(b) non-i.i.d.
Fig. 5: Test accuracy versus the number of time slots under different transmission strategies on the MNIST dataset: (a) i.i.d. data distribution and (b) non-i.i.d. data distribution.
@@ VII-D Comparison With Benchmark Schemes
    $\underset{p_{n,t}}{\text{min}}{}{}$ $\frac{2VLG^{2}q_{n,t}(h_{n,t},p_{n,t})}{N\sqrt{KT}}+Q_{n,t}e_{n,t}$ (57)
@@
    where $V>0$ is a tradeoff hyperparameter, and $Q_{n,t}$ is the virtual queue of device $n$. The queuing dynamics of $Q_{n,t}$ is given by $Q_{n,t+1}=Q_{n,t}+e_{n,t}-u_{n,t}.$
In Fig. 5, we compare the learning performance of the proposed algorithms with benchmark schemes on the MNIST dataset under both i.i.d. and non-i.i.d. data distributions. As observed, the proposed algorithms achieve near-optimal learning performance and outperform other benchmarks. The greedy optimality gap minimization scheme myopically uses up all remaining energy at each time slot, leading to poor performance in subsequent time slots. The Lyapunov-based online optimization scheme does not utilize the statistical knowledge of system dynamics, e.g., the distributions of the channel gains and the amount of harvested energy, resulting in poor long-term performance. In contrast, the proposed algorithms leverage the previously known or learned knowledge of channel and energy dynamics to design the transmission policy from a long-term perspective, thereby improving the learning performance.
Fig. 6 compares the proposed algorithms with benchmark methods on the CIFAR-10 dataset. To illustrate the broad applicability of the proposed methods, we implement CNNs as the classifier models, which do not satisfy Assumption 1. Despite this relaxation, the empirical results exhibit consistent performance trends, with the proposed algorithms significantly outperforming other benchmarks and closely approaching the optimal benchmark. This highlights the robustness and effectiveness of the proposed approaches under various FL scenarios.